OpenStudy (anonymous):

Assume that f(0) = 2 and f '(0) = 3. Calculate the derivative of the following function at x = 0...... (f(x))^3

myininaya (myininaya):

of the following function?

myininaya (myininaya):

because if we are just computing the derivative of the function f at x=0 then that is already given

OpenStudy (anonymous):

Sorry! Forgot the final bit

myininaya (myininaya):

USe the chain rule there

myininaya (myininaya):

to differentiate (f(x))^3

OpenStudy (anonymous):

That's where I'm at a loss. I get 3f'(x) and I thought that'd just be 9 but it's not...

myininaya (myininaya):

ok well do you the chain rule is this: \[[f(g(x))]'=g'(x) \cdot f'(g(x))\]

OpenStudy (anonymous):

oh nvm I forgot I was using an exponent. Time to redo everything!

OpenStudy (anonymous):

Thanks though

myininaya (myininaya):

let me change those function names up a little \[[g(f(x))]'=f'(x) \cdot g'(f(x))\]

myininaya (myininaya):

using this bottom thing I said here g=x^3 do you know how to different g?

OpenStudy (anonymous):

Indeed I do. I just realized I did the wrong thing and didn't use the exponent rule

myininaya (myininaya):

\[[g(f(x))]'=f'(x) \cdot g'(f(x)) =f'(x) \cdot 3(f(x))^2\] oh okay! so you got this then?

OpenStudy (anonymous):

Yes! Thanks a lot! Sometimes it just takes a fresh brain to help see what I did wrong.